OpenStudy (anonymous):

convert 89/6 to a mixed number

zepdrix (zepdrix):

`One full unit` can be expressed as \(\dfrac{6}{6}\). `Two full units` can be expressed as \(\dfrac{12}{6}\) We have \(-\dfrac{89}{6}\). We want to see how many full units we can pull out of this fraction. If you punch this fraction into your calculator \(89 \div 6\), you get a value of \(14.8333333\). Which is telling us that this fraction can be written as \(14\) whole units with some remainder left over. So let's multiply \(6\) and \(14\) and see how much we're left over with. \(6\cdot14=84\). We'll subtract this amount from our 89. \(89-84=5\). So we can write this fraction as \(14\) whole units, and \(5\) out of \(6\) left over. \[\large -\frac{89}{6} \quad = \quad -14\frac{5}{6}\]

zepdrix (zepdrix):

Oh sorry, there originally was a negative sign when you posted the question :) Just ignore the negatives in my comments then.